the ratio of the measures of two supplementary angles is 7:2 . find the value of x and the measures of each angle

Answer by edjones(8007)

Let x and y be the 2 angles.
x+y=180
7x=2y
x=2y/7
2y/7 + y = 180
2y+7y=1260
9y=1260
y=140
x=280/7=40
.
Ed